OpenStudy (anonymous):

how do you simplify: (4x√9x)-(6√x^3)

OpenStudy (anonymous):

\[4x\sqrt{9x}-6\sqrt{x^3}=4x\times 3\sqrt{x}-6x\sqrt{x}\] \[=12x\sqrt{x}-6x\sqrt{x}=6x\sqrt{x}\]